Around 13:30 UTC we put down INJ_MAIN and NARM/WARM nodes, opened BPC (DC on) to allow the work in the laser lab by the Artemis crew.
Around 18:10 LT we went on the IB platform to examine the ground connection of the PSTAB box installed on SIB1.
We found the west flange with the fedd-through of the PSTAB cables as indicated by Jean-Pierre Coulon (see picture 1). The small black wire starting from the PSTAB connector is attached to the metallic structure by means of a crocodile connector (see picture 2,3). We measured the resistance by means of an ohmmeter between:
1) exposed metallic plug before the crocodile connection and the tower = 0.3 Ohm after a few seconds (starting at 1.4 Ohm)
2) attached chassis and the tower = 0.1 Ohm
3) the metallic cylinder holding the cables in place and the tower = 3.4 Ohm
(see picture 4 as a reference).
After we went back to the control room, we closed again BPC loop, resumed INJ_MAIN, sent IMC_RESTORED as a request and then were able to relock the arms with the IR by using NARM and WARM metatron nodes. We leave the ITF like this.